# Virtual Target Trajectory Prediction for Stochastic Targets

### v1

@@ 6.1.1 Stochastic Maneuvers
Figure 10: Learned probability density function of the position of the target at different times. Figure reproduced from [18].
6.1.2 Deterministic Maneuvers
[…]
(a) 0 s
(b) 45 s
(c) 90 s
Figure 11: Test data: Histograms of samples of the probability density function of the position of the target at different times. Figure reproduced from [18].
@@ 6.1.2 Deterministic Maneuvers
Figure 12: Learned distribution: Histograms of samples of the probability density function of the position of the target at different times which does not change maneuvers. Figure reproduced from [18].
With the help of the new training data, the model is trained to predict the PDF of the position of the target at different times. Figure 12 depicts samples of the learned distribution of the predicted target positions for different times. When comparing it to the test data in Fig. 12, it can be seen that the model is able to predict the position of the target quite well. Nevertheless, there are some outliers, which are caused by the fact that the values of the learned PDF are not exactly zero. Thus, the model predicts a very low, but non-zero, probability for the target to be at a position other than the three possible positions.
@@ 6.1.3 Ballistic Targets
Figure 14 depicts the samples of $p(\mathbf{x}\mid t,\psi)$ created for $t\in\{0,45,90\}$ s into the flight of the ballistic trajectory and $\psi=\mathrm{BC}=500~{}\mathrm{kg/m}^{2}$. When comparing the learned distribution to the test data depicted in Fig. 14, which was obtained through costly Monte Carlo simulation, we observe that the model is able to predict the distribution of the ballistic trajectory quite well. While there are some samples (e.g., three samples in Fig. 13(f)) that do not fit the training data, the overall shape of the distribution is similar to the test data depicted in Fig. 14, with a correct position of the mean of the distribution and a correct shape of the distribution. Comparing this to the total of $10^{4}$ samples generated by the model, the number of such anomalies is considrabily low.
@@
Figure 13: Monte Carlo simulation: samples of the probability density function of the position of the target with $\mathrm{BC}=500~{}\frac{\mathrm{kg}}{\mathrm{m}^{2}}$ at different times. Figure reproduced from [18].
(d) 0 s
(e) 45 s
(f) 90 s
Figure 14: Samples of the learned probability density function of the position of the target with the additional parameter $\psi=\mathrm{BC}=500~{}\mathrm{kg/m}^{2}$ at different times. Figure reproduced from [18].
@@ 6.2 Generation of Virtual Target Trajectories
Number of samples per target and time step $10^{4}$
@@ 6.2.1 Single Target
Figure 15: Prediction of virtual target trajectories for a single real target (initial position: $[0,0]~{}\mathrm{km}$, flying northbound) with (a) one, (b) two, (c) three, and (d) four virtual targets and the respective clustered samples at time 90 s.
@@ 6.2.2 Multiple Targets
Figure 16 shows the prediction for three real spatially separated targets with one to four virtual targets. When only one virtual target is predicted for each real target, the virtual target trajectories are similar to the average of the real target trajectories (Fig. 16(a)). Increasing the number of virtual targets to two or three still results in the prediction of straight trajectories, just the position of the virtual targets is shifted (Figures 16(b)16(c)). When the number of virtual targets is higher than the number of real targets (Fig. 16(d)), the samples of one real target are clustered into multiple virtual targets, similar to Fig. 15.
When the real targets are close to each other, as shown in Fig. 17, the distributions of the target positions overlap, leading to one left turn, one right turn, and one straight flight as virtual target trajectories.
@@
Figure 16: Prediction of virtual target trajectories for three real spatially separated targets (initial positions: $[0,-40]~{}\mathrm{km},[0,0]~{}\mathrm{km},[0,40]~{}\mathrm{km}$, all targets flying northbound) with (a) one, (b) two, (c) three, and (d) four virtual targets and the respective clustered samples at time 90 s.
@@
Figure 17: Prediction of virtual target trajectories for three real targets close to each other (initial positions: $[0,-5],~{}\mathrm{km}[0,0]~{}\mathrm{km},[0,5]~{}\mathrm{km}$, all targets flying northbound) with (a) one, (b) two, (c) three, and (d) four virtual targets and the respective clustered samples at time 90 s.
@@ 7 Conclusion
Using Conditional Normalizing Flows, the position of the target at any given time can be predicted either by directly calculating the learned probability density function or by sampling from it. This approach is useful in various applications where predicting the target’s position is essential. Compared to Monte Carlo simulations, the approach is computationally more efficient and allows for the calculation of the probability density function of the target’s position. Since most targets do not follow perfectly deterministic trajectories, the usage of a stochastic predictor can take the uncertainty of the target trajectory into account, potentially leading to more robust downstream applications. When deterministic trajectories are required, the samples can be clustered to obtain a set of representative trajectories. Thus, the presented approach allows to predict deterministic trajectories for stochastically moving targets, making it a versatile tool for trajectory prediction. It can be used as a drop-in replacement for deterministic trajectory predictions used in areas such as guidance laws, path planning, and other applications where a prediction of the target’s trajectory is needed.

### v2

@@ 6.1.2 Deterministic Maneuvers
Figure 12: Test data: Histograms of samples of the probability density function of the position of the target at different times. Figure reproduced from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).
@@
Figure 13: Learned distribution: Histograms of samples of the probability density function of the position of the target at different times. Figure reproduced from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).
With the help of the new training data, the model is trained to predict the PDF of the position of the target at different times. Figure 13 (reproduced from [20]) depicts samples of the learned distribution of the predicted target positions for different times. When comparing it to the test data in Fig. 13, it can be seen that the model is able to predict the position of the target quite well. Nevertheless, there are some outliers, which are caused by the fact that the values of the learned PDF are not exactly zero. Thus, the model predicts a very low, but non-zero, probability for the target to be at a position other than the three possible positions.
@@ 6.1.3 Ballistic Targets
Figure 14 (reproduced from [20]) depicts the samples of $p(\mathbf{x}\mid t,\psi)$ created for $t\in\{0,45,90\}$ s into the flight of the ballistic trajectory and $\psi=\xi=500~\mathrm{kg/m}^{2}$. When comparing the learned distribution to the test data depicted in Fig. 9, which was obtained through costly Monte Carlo simulation, we observe that the model is able to predict the distribution of the ballistic trajectory quite well. While there are some samples (e.g., three samples in Fig. 14(c)) that do not fit the training data, the overall shape of the distribution is similar to the test data depicted in Fig. 9, with a correct position of the mean of the distribution and a correct shape of the distribution. Comparing this to the total of $10^{4}$ samples generated by the model, the number of such anomalies is considrabily low.
@@
Figure 14: Samples of the learned probability density function of the position of the target with the additional parameter $\psi=\xi=500~\mathrm{kg/m}^{2}$ at different times. Figure reproduced from Schneider, M., Loureiro, R., Cunis, T., and Fichter, W., “Trajectory Prediction for Missile Targets: A Probabilistic Approach Using Machine Learning,” in CEAS EuroGNC Conference, 2024. Licensed under a Creative Commons Attribution 4.0 International License ($C$-BY 4.0).
@@ 6.2 Generation of Virtual Target Trajectories
Number of samples per target and time step 200200
@@ 6.2.1 Single Target
Figure 15: Prediction of virtual target trajectories for a single real target (initial position: $0,0]~\mathrm{km}$, flying northbound) and the respective clustered samples at 90 s.
@@ 6.2.2 Multiple Targets
Figure 16 shows the prediction for three real spatially separated targets with one to four virtual targets. When only one virtual target is predicted for each real target, the virtual target trajectories are similar to the average of the real target trajectories (Fig. 16(a)). Increasing the number of virtual targets to two or three still results in the prediction of straight trajectories, just the position of the virtual targets is shifted (Figures 16(b)16(c)). When the number of virtual targets is higher than the number of real targets (Fig. 16(d)), the samples of one real target are clustered into multiple virtual targets, similar to Fig. 15. When the real targets are close to each other, as shown in Fig. 17, the distributions of the target positions overlap, leading to one left turn, one right turn, and one straight flight as virtual target trajectories.
@@
Figure 16: Prediction of virtual target trajectories for three real spatially separated targets (initial positions: $0,-40]~\mathrm{km},[0,0]~\mathrm{km},[0,40]~\mathrm{km}$, flying northbound) and the respective clustered samples at 90 s.
@@
Figure 17: Prediction of virtual target trajectories for three real targets close to each other (initial positions: $0,-5]~\mathrm{km},[0,0]~\mathrm{km},[0,5]~\mathrm{km}$, flying northbound) and the respective clustered samples at 90 s.
@@ 7 Conclusion
Using Conditional Normalizing Flows, the position of the target at any given time can be predicted either by directly calculating the learned probability density function or by sampling from it. This approach is useful in various applications where predicting the target’s position is essential. Compared to Monte Carlo simulations, the approach can be computationally more efficient while also providing exact density evaluation: for the simple maneuver scenario creating $10^{4}$ samples the CNF required about 0.17 s compared to 6.64 s for the Monte Carlo reference, and for the ballistic scenario 0.27 s versus 178 s for Monte Carlo (see Sec. 6.1). These timing measurements illustrate the potential computational advantage of the learned model for large-sample evaluations. Since most targets do not follow perfectly deterministic trajectories, the usage of a stochastic predictor can take the uncertainty of the target trajectory into account, potentially leading to more robust downstream applications. When deterministic trajectories are required, the samples can be clustered to obtain a set of representative trajectories. Thus, the presented approach allows to predict deterministic trajectories for stochastically moving targets, making it a versatile tool for trajectory prediction. It can be used as a drop-in replacement for deterministic trajectory predictions used in areas such as guidance laws, path planning, and other applications where a prediction of the target’s trajectory is needed.